OpenStudy (anonymous):

Find the equation of the parabola: y=ax^2+bx+c that passes through (2,0), (3,-1), and (4,0).

OpenStudy (anonymous):

You need to solve the 3 equations below \[ 4 a+2 b+c=0\\ 9 a+3 b+c=-1\\ 16 a+4 b+c=0 \]

OpenStudy (anonymous):

You get \[\begin{array}{ccc} a=1; & b=-6; & c=8; \end{array} \]

OpenStudy (anonymous):

Ohhhh. I'm stupid. Okay thanks

OpenStudy (anonymous):

yw